# subfigure0(#1)

### v2

@@ 3.2. Fruchterman-Reingold for ball graphs
Classical Fruchterman-Reingold algorithms come in two types: at each iteration, either all vertices move at once according to their displacement vectors, or the vertices move one at a time, requiring a recalculation of forces at each iteration. In the former case, the new position for each vertex $v$ is $\left[\begin{matrix}x_{v}&y_{v}\end{matrix}\right]^{\mathsf{T}}+\delta_{v}$. In the latter case, we cycle through the vertices of $G$, moving a single vertex $v$ according to $\delta_{v}$ at each iteration. Our solution to the problem of breaking edges is essentially the same in either case: if any pair of vertices previously within distance $1$ would be at distance $>1$ after any iteration, we avoid moving any vertices at all, lower the temperature by $t/(\mathrm{it}+1)$, where ‘$\mathrm{it}$’ denotes the total number of iterations, and move on to the next iteration (essentially weakening the forces $F_{u,v}$). If all pairs of vertices previously within distance $1$ remain within distance $1$, we keep the new positions and again lower the temperature by $t/(\mathrm{it}+1)$ for the next iteration. As the temperature decreases to $t/(\mathrm{it}+1)$ at the last iteration, the positions of the vertices stabilize.
4. Formation planning for autonomous swarms
In this section, we apply Algorithm 1 and the modified Fruchterman-Reingold algorithm from Section 3.2 to obtain swarm formations with even area coverage and high reliability. Returning to our running example, we consider an autonomous swarm of satellites assigned to image a region on the surface of an object. We begin by assigning satellites to image the outer boundary of the region. Our goal is to assign the remaining satellites to cover the interior of the region so that the entire region is evenly covered and so that the communication network is not only connected but highly reliable. We note that search and rescue missions by robots, imaging of an ocean bed by autonomous underwater vehicles, or use of unmanned aerial vehicles for precision agriculture may be modeled similarly. Our findings indicate that using Algorithm 1 alone (as opposed to using it in conjunction with our modified Fruchterman-Reingold algorithm) is particularly effective in accomplishing this goal.
Figure 3. Fifteen agents cover the outer boundary of a circular region, modeled by a unit disk graph forming a regular $15$-gon whose edges are of length $0.9$
@@ 4. Formation planning for autonomous swarms
    Random & Spring: Add vertices as in item (M1), but apply the Fruchterman-Reingold algorithm from Section 3.2 each time a vertex is added, leaving the positions of the initial $15$ vertices fixed.
@@
    Buffer & Spring: Add vertices as in item (M3), but apply the Fruchterman-Reingold algorithm from Section 3.2 each time a vertex is added.
The specifics of each of these methods are discussed below, and examples of the resulting graphs are depicted in Figure 4. For a reliability measure, we use the all-terminal reliability with edge-operation probabilities all $90\%$. These edge-operation probabilities may be adjusted, and unreliable vertices introduced, with additional information regarding the agents in the application and their communication devices. Using the Zero-One Estimator theorem of Karp and Luby [25], or one of its improved forms (see, e.g.[14]), one can easily compute upper bounds on the number of Monte Carlo simulations necessary to obtain, with probability $1-\delta$, an estimate of $\operatorname{Rel}_{A}$ with relative error $\varepsilon$ (known as an $(\varepsilon,\delta)$-approximation). For a graph $G$ with high reliability, however, these upper bounds can be quite large. Such Monte Carlo techniques proved inefficient for the more reliable graphs produced by the methods in items (M3) and (M4) above. Instead, we use an open-source program called $\mathcal{K}$-RelNet, introduced in [30], to obtain $(\varepsilon,\delta)$-estimates of $\operatorname{Rel}_{A}(G)$. Essentially, the program reduces the problem of finding all-terminal reliability (or the more general $K$-terminal reliability, see [12]) to counting the number of assignments satisfying a Boolean formula in conjunctive normal form. It then obtains an $(\varepsilon,\delta)$-approximation of this count using a state-of-the-art approximate model counter ApproxMC, introduced in [10] and now in its sixth version, ApproxMC6 [37]. We set $\varepsilon=0.8$ and $\delta=0.2$ for our relative estimates; these are the default values for the $\mathcal{K}$-RelNet program, and a relative error of $0.8$ provides quite a small absolute error bound for the more reliable graphs in question.
@@
Figure 4. Fifteen vertices are added to the interior of a region outlined by a regular $15$-gon with edge-lengths of $0.9$ according to methods (M1)-(M4).
As a measure of area coverage, we calculate the radius of the largest empty circle (containing no vertices of the graph) whose center is contained in the convex hull of the vertex set. The region will be well-covered if the radius of the largest empty circle is small. A largest empty circle whose center is contained in the convex hull of a set of $n$ points can be found in $\mathcal{O}(n\log{n})$ time [34]. In Table 1, we record the average all-terminal reliability and the average radius of the largest empty circle over $100$ graphs obtained using each of the methods (M1)-(M4). The results in Table 1 are compared, and standard deviations depicted, in Figure 5.
@@
Table 1. The average all-terminal reliability $\operatorname{Rel}_{A}$ with edge-operation probabilities of $90\%$ and the average radius of the largest empty circle (LEC) over $100$ graphs obtained using each method (M1)-(M4).
@@
Figure 5. A comparison of average all-terminal reliability (with edge-operation probabilities of $90\%$) of graphs obtained via methods (M1)-(M4). Shaded regions depict one standard deviation.
@@
The Fruchterman-Reingold algorithm has quite a different effect when applied to the graphs obtained via method (M1) than one might expect: vertices are pushed away from the interior of the region towards the boundary. This increases, rather than decreases, the uncovered area of the region. On the other hand, it also has the effect of slightly increasing the reliability. We compared both types of our modified Fruchterman-Reingold algorithm from Section 3.2 (either all non-fixed vertices move at each iteration, or vertices move one at a time), and found that the former method was more effective for both reliability and area coverage purposes.
Averaged over $100$ trials, the all-terminal reliability of the graph obtained by adding $15$ vertices one at a time uniformly at random to the interior of the region, applying our modified Fruchterman-Reingold algorithm after each added vertex, is about $85.16\%$. The average radius of the largest empty circle is about $1.43$. Figure 5(a) plots the average reliabilities obtained using method (M1) against those using method (M2).
@@
We average the all-terminal reliability of $100$ graphs obtained by adding $15$ vertices to the interior of a region outlined by $15$ satellites using Algorithm 1 ($b=0.65$) and choosing the most reliable candidate at each step. The average reliability after adding $15$ vertices is about $99.09\%$. The average radius of the largest empty circle is about $0.67$. The results are displayed in Figure 5(b).
Finally, we consider method (M4). Each time we add a vertex to the interior of the region to maximize reliability while being of distance at least $0.65$ from every other vertex, we apply our Fruchterman-Reingold algorithm from Section 3.2. We again use the version of our algorithm in which all non-fixed vertices move at each iteration. Note that this algorithm may move vertices which were placed outside of buffer distance closer together. Indeed, the largest empty circle is, on average, larger in this case. Over $100$ trials, the average radius of the largest empty circle is about $0.97$. The average all-terminal reliability after adding $15$ vertices is about $99.2\%$. The results are displayed in Figure 5(c).
5. Future directions
@@
In a similar vein, we have not fully resolved the problem of finding the most reliable unit ball graphs of a given order which cover a given area (even for the circular area considered in Section 4). This problem is likely a difficult one, as finding most reliable (abstract) graphs of a given order and size remains an open problem. However, the constraints here are quite different, and we pose this problem for completeness.
@@ Appendix A Vertices on the boundary
Let $V\subset\mathbb{R}^{2}$ be a finite set of points in general position, and let $\emptyset\neq S\subseteq V$. If $S$ is contained in a $(b,c)$-annulus with all points in $V-S$ outside of its outer boundary, then there exists a second $(b,c)$-annulus with two points $u,v\in V$ on its boundary which contains $S-\{u,v\}$ and is such that all points in $V-(S\cup\{u,v\})$ lie outside of the outer boundary.
@@
First, let $A$ be a $(b,c)$-annulus with $S\subset A$ and $V-S$ lying outside of the outer boundary of $A$. It is not hard to see that we can shift $A$ slightly so that some point $u\in V$ is on its outer or inner boundary. We now rotate this shifted annulus around the fixed point $u$ until a second point $v$ lies on its boundary to obtain the desired annulus.
@@ Appendix B InitialSet algorithm
Acknowledgments
This work was funded in part by NASA under cooperative agreement VT-80NSSC20M0213. C.B. acknowledges support from the Vermont Space Grant Consortium Graduate Fellowship Program.

### v3

@@ III-B Fruchterman-Reingold for ball graphs
Classical Fruchterman-Reingold algorithms come in two types: at each iteration, either all vertices move at once according to their displacement vectors, or the vertices move one at a time, requiring a recalculation of forces at each iteration. In the former case, the new position for each vertex $v$ is $\left[\begin{matrix}x_{v}&y_{v}\end{matrix}\right]^{\mathsf{T}}+\delta_{v}$. In the latter case, we cycle through the vertices of $G$, moving a single vertex $v$ according to $\delta_{v}$ at each iteration. Our solution to the problem of breaking edges is essentially the same in either case: if any pair of vertices previously within distance $1$ would be at distance >1>1 after any iteration, we avoid moving any vertices at all, lower the temperature by $t/(\mathrm{it}+1)$, where $\mathrm{it}$’ denotes the total number of iterations, and move on to the next iteration (essentially weakening the forces $F_{u,v}$). If all pairs of vertices previously within distance $1$ remain within distance $1$, we keep the new positions and again lower the temperature by $t/(\mathrm{it}+1)$ for the next iteration. As the temperature decreases to $t/(\mathrm{it}+1)$ at the last iteration, the positions of the vertices stabilize.
IV Formation planning for autonomous swarms
In this section, we apply Algorithm 1 and the modified Fruchterman-Reingold algorithm from Section III-B to obtain swarm formations with even area coverage and high reliability. Returning to our running example, we consider an autonomous swarm of satellites assigned to image a region on the surface of an object. We begin by assigning satellites to image the outer boundary of the region. Our goal is to assign the remaining satellites to cover the interior of the region so that the entire region is evenly covered and so that the communication network is not only connected but highly reliable. We note that search and rescue missions by robots, imaging of an ocean bed by autonomous underwater vehicles, or use of unmanned aerial vehicles for precision agriculture may be modeled similarly. Our findings indicate that using Algorithm 1 alone (as opposed to using it in conjunction with our modified Fruchterman-Reingold algorithm) is particularly effective in accomplishing this goal.
Figure 3: Fifteen agents cover the outer boundary of a circular region, modeled by a unit disk graph forming a regular $15$-gon whose edges are of length $0.9$
@@ IV Formation planning for autonomous swarms
    Random & Spring: Add vertices as in item (M1), but apply the Fruchterman-Reingold algorithm from Section III-B each time a vertex is added, leaving the positions of the initial $15$ vertices fixed.
@@
    Buffer & Spring: Add vertices as in item (M3), but apply the Fruchterman-Reingold algorithm from Section III-B each time a vertex is added.
The specifics of each of these methods are discussed below, and examples of the resulting graphs are depicted in Figure 4. For a reliability measure, we use the all-terminal reliability with edge-operation probabilities all $90\%$. These edge-operation probabilities may be adjusted, and unreliable vertices introduced, with additional information regarding the agents in the application and their communication devices. Using the Zero-One Estimator theorem of Karp and Luby [33], or one of its improved forms (see, e.g.[18]), one can easily compute upper bounds on the number of Monte Carlo simulations necessary to obtain, with probability $1-\delta$, an estimate of $\operatorname{Rel}_{A}$ with relative error $\epsilon$ (known as an $(\epsilon,\delta)$-approximation). For a graph $G$ with high reliability, however, these upper bounds can be quite large. Such Monte Carlo techniques proved inefficient for the more reliable graphs produced by the methods in items (M3) and (M4) above. Instead, we use an open-source program called $\mathcal{K}$-RelNet, introduced in [34], to obtain $(\epsilon,\delta)$-estimates of $\operatorname{Rel}_{A}(G)$. Essentially, the program reduces the problem of finding all-terminal reliability (or the more general $K$-terminal reliability, see [16]) to counting the number of assignments satisfying a Boolean formula in conjunctive normal form. It then obtains an $(\epsilon,\delta)$-approximation of this count using a state-of-the-art approximate model counter ApproxMC, introduced in [35] and now in its sixth version, ApproxMC6 [36]. We set $\epsilon=0.8$ and $\delta=0.2$ for our relative estimates; these are the default values for the $\mathcal{K}$-RelNet program, and a relative error of $0.8$ provides quite a small absolute error bound for the more reliable graphs in question.
@@
Figure 4: Fifteen vertices are added to the interior of a region outlined by a regular $15$-gon with edge-lengths of $0.9$ according to methods (M1)-(M4).
As a measure of area coverage, we calculate the radius of the largest empty circle (containing no vertices of the graph) whose center is contained in the convex hull of the vertex set. The region will be well-covered if the radius of the largest empty circle is small. A largest empty circle whose center is contained in the convex hull of a set of $n$ points can be found in $\mathcal{O}(n\log{n})$ time [37]. In Table I, we record the average all-terminal reliability and the average radius of the largest empty circle over $100$ graphs obtained using each of the methods (M1)-(M4). The results in Table I are compared, and standard deviations depicted, in Figure 5.
TABLE I: The average all-terminal reliability $\operatorname{Rel}_{A}$ with edge-operation probabilities of $90\%$ and the average radius of the largest empty circle (LEC) over $100$ graphs obtained using each method (M1)-(M4).
@@
Figure 5: A comparison of average all-terminal reliability (with edge-operation probabilities of $90\%$) of graphs obtained via methods (M1)-(M4). Shaded regions depict one standard deviation.
@@
The Fruchterman-Reingold algorithm has quite a different effect when applied to the graphs obtained via method (M1) than one might expect: vertices are pushed away from the interior of the region towards the boundary. This increases, rather than decreases, the uncovered area of the region. On the other hand, it also has the effect of slightly increasing the reliability. We compared both types of our modified Fruchterman-Reingold algorithm from Section III-B (either all non-fixed vertices move at each iteration, or vertices move one at a time), and found that the former method was more effective for both reliability and area coverage purposes.
Averaged over $100$ trials, the all-terminal reliability of the graph obtained by adding $15$ vertices one at a time uniformly at random to the interior of the region, applying our modified Fruchterman-Reingold algorithm after each added vertex, is about $85.16\%$. The average radius of the largest empty circle is about $1.43$. Figure 5a plots the average reliabilities obtained using method (M1) against those using method (M2).
@@
We average the all-terminal reliability of $100$ graphs obtained by adding $15$ vertices to the interior of a region outlined by $15$ satellites using Algorithm 1 ($b=0.65$) and choosing the most reliable candidate at each step. The average reliability after adding $15$ vertices is about $99.09\%$. The average radius of the largest empty circle is about $0.67$. The results are displayed in Figure 5b.
Finally, we consider method (M4). Each time we add a vertex to the interior of the region to maximize reliability while being of distance at least $0.65$ from every other vertex, we apply our Fruchterman-Reingold algorithm from Section III-B. We again use the version of our algorithm in which all non-fixed vertices move at each iteration. Note that this algorithm may move vertices which were placed outside of buffer distance closer together. Indeed, the largest empty circle is, on average, larger in this case. Over $100$ trials, the average radius of the largest empty circle is about $0.97$. The average all-terminal reliability after adding $15$ vertices is about $99.2\%$. The results are displayed in Figure 5c.
V Future directions
@@
In a similar vein, we have not fully resolved the problem of finding the most reliable unit ball graphs of a given order which cover a given area (even for the circular area considered in Section IV). This problem is likely a difficult one, as finding most reliable (abstract) graphs of a given order and size remains an open problem. However, the constraints here are quite different, and we pose this problem for completeness.
@@ Appendix A Vertices on the boundary
Let $V\subset\mathbb{R}^{2}$ be a finite set of points in general position, and let $\emptyset\neq S\subseteq V$. If $S$ is contained in a $(b,c)$-annulus with all points in VSV-S outside of its outer boundary, then there exists a second $(b,c)$-annulus with two points $u,v\in V$ on its boundary which contains $S-\{u,v\}$ and is such that all points in $V-(S\cup\{u,v\})$ lie outside of the outer boundary.
@@
First, let $A$ be a $(b,c)$-annulus with $S\subset A$ and VSV-S lying outside of the outer boundary of $A$. It is not hard to see that we can shift $A$ slightly so that some point $u\in V$ is on its outer or inner boundary. We now rotate this shifted annulus around the fixed point $u$ until a second point $v$ lies on its boundary to obtain the desired annulus.